# Self-Employment as a Signal: Career Concerns with Hidden Firm Performance

### v1

@@ 7.1 Worker learning about own type
The absorbing logic extends: if $E$ is optimal at $(\mu_{t},m_{t})$, then the public state remains $(\alpha,\beta)$ and the same comparison repeats. [^19] Private updating under $E$ can move $m_{t+1}$, but with flat wages independent of private realizations, there is a neighborhood of private beliefs for which absorption persists. Public learning is restored only if the market sees an informative signal from $E$. Private learning under $E$ may move $m_{t+1}$ but does not overturn absorption locally with flat wages. Under equilibrium posting (application inference), the cutoff $\hat{m}(\alpha,\beta)$ shifts down due to the adverse-selection discount.
@@ 7.2 Partial public observability inside firms
Suppose that under $E$ the market observes a coarse signal $z_{t}\in\{0,1\}$ with informativeness $\phi\in[0,1)$:
$\Pr(z_{t}=1\mid\theta,E)=\phi\,\theta+(1-\phi)\,\bar{z}.$
For sufficiently small $\phi$, the equilibrium still features an absorbing hidden-employment region; as $\phi$ increases, the region shrinks and vanishes only when $E$ generates near-parity information with $S$.
@@ 7.3 Simple screening menus
Allow firms to post menus $(w,p)$ where $p$ is the intensity of public-facing tasks (or disclosure). Under competitive posting and zero profits, menus cannot separate types beyond the entry cutoff: raising $p$ to attract higher types requires lowering $w$, which erodes the insurance value and pushes those types back to $S$. The main margin remains entry vs. self-employment; menus tilt the cutoff but do not overturn absorption.
@@ 8 Conclusion
Three forces organize the results. First, an insurance–information trade-off: hidden employment offers insurance but suppresses public learning and future outside options. Second, selection: relatively stronger types opt to produce public signals, making observed outcomes more informative. Third, inference from applications: equilibrium wage posting discounts for adverse selection in the applicant pool. Together these forces generate an absorbing hidden-employment region—a set of public beliefs for which workers optimally accept hidden employment forever—alongside switching back to public production when reputation is sufficiently strong.
We deliver a simple characterization under both posting regimes. In the benchmark (myopic) case, wages equal the public posterior mean and the absorbing region expands with risk aversion and with career length (as single outcomes move beliefs less). Under equilibrium posting, a fixed point links the acceptance cutoff to the truncated-posterior wage; existence and uniqueness follow from monotonicity of the truncated mean and single crossing in the worker’s problem. Pricing sophistication compresses wages relative to the benchmark at a given public state and tightens self-selection, shifting the cutoff down.
The framework yields transparent empirical implications. Early in careers, when public signals move beliefs more, workers are more willing to generate public information; later, absorbing hidden employment becomes more prevalent. After public successes (failures), the benchmark predicts wide wage dispersion, whereas equilibrium posting corrects for winner’s-curse effects through application-based inference, compressing the success–failure gap. Across occupations, environments with portable public histories (e.g., open portfolios, platform ratings, contest leaderboards) should feature stronger selection into public production by higher-ability workers and sharper wage responses to public outcomes. [^20] Examples include code and data science (open repositories), academic research (working papers and preprints), design (public portfolios), and online freelancing (platform ratings).
@@
Two design lessons follow. First, organizational transparency is a lever: policies that selectively externalize performance (client-visible tasks, standardized public metrics) can substitute for self-employment as a reputational technology, [^21] For instance, making some client KPIs or review summaries public, adopting external code review, or giving structured, verifiable titles that map to outside-market signals. reducing the absorbing hidden-employment region and speeding market learning. Second, marketplace curation (platform labels, verified histories) changes the relative returns to public signals and therefore the equilibrium mix of public versus hidden production.
@@
By tying a classic career-concerns logic to environments with coexisting public and hidden performance, the paper provides a compact theory of when employment becomes absorbing, how application-based inference reshapes wages, and why the returns to public production are largest early in careers. We hope this framework proves useful for organizing empirical work on modern reputation systems and for guiding organizational choices about transparency and the design of early-career opportunities.
Acknowledgements
We thank Emiliano Catonini and Steven Kivinen for very helpful comments and encouragement on earlier drafts and related work. We are grateful to seminar and workshop participants at our home institutions and elsewhere for suggestions that improved the paper. All remaining errors are our own.
Appendix A: Proofs
@@ Acknowledgements
If $\theta\leq\hat{\theta}(\alpha,\beta)$ then $\Delta(\theta;\alpha,\beta,w(\alpha,\beta))\leq 0$. Employment leaves $(\alpha,\beta)$ (and hence $w(\alpha,\beta)$) unchanged, so the same inequality holds next period. By induction, employment is repeatedly optimal. ∎

### v3

@@ 6.1 Worker learning about own type
The absorbing logic extends: if $E$ is optimal at $(\mu_{t},m_{t})$, then the public state remains $(\alpha,\beta)$ and the same comparison repeats. [^19] Private updating under $E$ can move $m_{t+1}$, but with flat wages independent of private realizations, there is a neighborhood of private beliefs for which absorption persists. Public learning is restored only if the market sees an informative signal from $E$. Private learning under $E$ may move $m_{t+1}$ but does not overturn absorption locally with flat wages. Under sophisticated posting (application inference), the cutoff $\hat{m}(\alpha,\beta)$ shifts down due to the adverse-selection discount.
@@ 6.2 Partial public observability inside firms
So far, employment at the firm $E$ has been fully opaque to the outside market: individual performance affects wages only through its effect on the worker’s continuation value inside the firm. In many settings, however, firm jobs generate some coarse public signals—for example, occasional client feedback, standardized performance ratings, or sporadic disclosure of individual outcomes. To capture this, suppose that under $E$ the market observes a binary signal $z_{t}\in\{0,1\}$ with informativeness $\phi\in[0,1]$:
$\Pr(z_{t}=1\mid\theta,E)\;=\;\phi\,\theta+(1-\phi)\,\bar{z},$
where $\bar{z}\in(0,1)$ is a baseline success probability independent of $\theta$. When $\phi=0$, the signal is uninformative about talent and the extension collapses to the baseline model in which $E$ generates no public information. As $\phi$ increases, $z_{t}$ becomes more informative about $\theta$ in the sense of Blackwell: a larger share of the signal is driven by the same $\theta$-dependent component that governs public output under $S$. In the limit, as $\phi\to 1$, $E$ and $S$ generate nearly parity information about talent.
This partial observability changes the worker’s dynamic trade-off in two ways. First, for high-$\theta$ types, more informative in-firm signals make $E$ more attractive: choosing $E$ no longer freezes the public record, so a talented worker can rely on occasional successes under $E$ to sustain or improve her reputation. Second, for low-$\theta$ types, more informative in-firm signals make $E$ less attractive: opacity is less effective as a shield and the benefit of retreating into hidden employment is reduced.
Consequently, the absorbing hidden-employment region—the set of public states $(\alpha,\beta)$ at which it is optimal for all types to choose $E$ forever and stop actively managing their public record—shrinks as $\phi$ increases. For sufficiently low values of $\phi$, the region remains non-empty: if in-firm signals are very noisy, hidden employment is still effectively opaque and the analysis from Sections 45 carries through. As $\phi$ approaches one, the region eventually disappears: once $E$ and $S$ generate almost equally informative public records, there is no public state at which all types strictly prefer to remain in $E$ forever purely for informational reasons.
The following result formalizes this comparative static.
Proposition 6.
Consider the extension with in-firm signals $z_{t}$ as above and denote by $\mathcal{A}(\phi)\subseteq\{(\alpha,\beta)\}$ the absorbing hidden-employment region under informativeness parameter $\phi$, i.e., the set of public states $(\alpha,\beta)$ at which it is optimal for all $\theta\in[0,1]$ to choose $E$ forever. Then:
  1. 1.
    At $\phi=0$, $\mathcal{A}(0)$ coincides with the absorbing region characterized in Proposition 1 and is non-empty.
  2. 2.
    For $\phi$ in a neighbourhood of zero, $\mathcal{A}(\phi)$ remains non-empty.
  3. 3.
    If at $\phi=1$ the public signals generated under $E$ and $S$ are observationally equivalent (i.e., induce the same conditional distribution of public histories for every $\theta$), then $\mathcal{A}(\phi)$ shrinks as $\phi$ increases and is empty for all $\phi$ sufficiently close to one.
Proof sketch.
At $\phi=0$ the extension reduces to the baseline model, so $\mathcal{A}(0)$ is non-empty by Proposition 1. For $\phi$ in a neighbourhood of zero, continuity of the worker’s value function and of best responses in the signal distribution implies that the set of states at which $E$ strictly dominates $S$ for all types deforms only slightly. Hence, $\mathcal{A}(\phi)$ remains non-empty for all $\phi\in[0,\phi_{0})$ for some $\phi_{0}>0$.
As $\phi$ increases, the signal under $E$ becomes more informative about $\theta$ in the sense of Blackwell, and the transition kernel for public beliefs under $E$ converges to that under $S$. Under the maintained assumption that at $\phi=1$ the distribution of public histories is the same under $E$ and $S$ for any fixed $\theta$, the only remaining difference between the two actions is their one-period payoff streams.
For any public state $(\alpha,\beta)$ there are then types $\theta$ who strictly prefer $S$ (e.g., high types for whom the additional risk under $S$ is rewarded) and types who weakly prefer $E$. Thus, there is no state at which all $\theta$ strictly prefer $E$ forever and $\mathcal{A}(1)$ is empty. Upper hemicontinuity of $\mathcal{A}(\cdot)$ in $\phi$ then implies that $\mathcal{A}(\phi)=\emptyset$ for all $\phi$ sufficiently close to one. ∎
Proposition 6 shows that the absorbing hidden-employment region is robust to modest in-firm transparency but is not a knife-edge artifact of complete opacity. As long as in-firm signals remain sufficiently noisy, there are public states at which all types optimally retreat into hidden employment and allow the public record to stagnate. As in-firm performance becomes clearer to outside observers, this region shrinks and eventually vanishes: opacity inside firms ceases to be a distinct reputational option, and the worker’s problem increasingly resembles choosing between two jobs that generate comparable public records.
This extension connects our mechanism to literatures on partial transparency and signal design inside organizations, where the informativeness of performance measures can be chosen or regulated, and highlights that modest increases in in-firm observability need not overturn the central absorbing-region logic.
@@ 6.3 Simple screening menus
So far, employment at the firm $E$ has been represented by a single contract: a flat wage and full opacity to the outside market. Many real jobs, however, combine a fixed wage with a choice over how much of the worker’s output is made visible outside the firm (e.g. more client-facing work, more external presentations, or more disclosure of individual performance metrics). To capture this in reduced form, suppose firms can post simple menus of the form $(w,p)$, where $w$ is the wage and $p\in[0,1]$ is the intensity of public-facing tasks or disclosure in $E$. A higher $p$ means that a larger fraction of the worker’s performance under $E$ is transformed into public signals (in the same spirit as the partial observability parameter $\phi$ in the previous subsection), so that $p=0$ corresponds to a fully opaque firm job and $p=1$ (to first order) to a job that generates almost the same public record as $S$.
For a given menu $(w,p)$ and public state $(\alpha,\beta)$, the worker’s expected utility under $E$ trades off three effects. First, higher $w$ raises current consumption and provides insurance against output risk. Second, higher $p$ makes employment more like self-employment from a reputational point of view: the public record continues to evolve under $E$, which is attractive for high-$\theta$ types and unattractive for low-$\theta$ types. Third, the option value of switching back to self-employment in the future depends on how the public state $(\alpha,\beta)$ evolves under $(w,p)$. For fixed $p$, there is single crossing in $\theta$ between $E$ and $S$ as in the baseline model: types above a cutoff prefer to bear the risk of $S$ in exchange for more informative public signals, while types below the cutoff prefer the firm’s insurance. For fixed $w$, there is also a natural single-crossing structure in $p$: high-$\theta$ types value being observed (higher $p$), low-$\theta$ types value opacity (lower $p$).
Under competitive posting and zero profits, wages must adjust to the menu’s informativeness. When a firm increases $p$ in order to make a contract more attractive to high types, the implied applicant pool becomes more positively selected, but the contract also exposes workers to more reputational risk. Zero profits then require lowering $w$ to keep expected output net of wages at zero. This has two implications. First, menus cannot generate rich separation within employment over and above the self-employment margin: any high-$p$ contract that succeeds in attracting high types must offer a wage low enough that some of those types are again better off choosing $S$, which undermines the intended screening. Second, the main margin remains entry versus self-employment: menus tilt the location of the cutoff $\hat{\theta}(\alpha,\beta)$ by making $E$ more or less attractive at the boundary, but they do not create additional disjoint regions in $(\alpha,\beta,\theta)$ space where different types sort into qualitatively different forms of employment inside the firm.
The upshot is that simple screening menus are unlikely to overturn the absorbing hidden-employment region. For sufficiently low $p$ (or for menus in which the high-$p$ option is not too attractive), the logic of Proposition 1 continues to apply: there remain public states $(\alpha,\beta)$ at which all types strictly prefer firm employment over self-employment, and once a worker enters that region and chooses $E$, the public record ceases to be actively managed. Allowing menus shifts the cutoff and thus the size of the absorbing region—for example, making transparent menus more widely available can shrink the region when workers are sufficiently high types—but does not eliminate the basic absorption logic. A full characterization of optimal menus and their interaction with dynamic reputation is beyond the scope of this paper; here the main lesson is that simple competitive screening devices at the firm level tilt the entry margin rather than provide a complete substitute for self-employment as a tool for reputation management.
6.4 Hidden effort and work–shirk dynamics
So far, self-employment affects only the observability of output: conditional on choosing $S$, success arrives with probability $\theta$ and failure with probability $1-\theta$. A natural extension is to allow the worker to choose a hidden effort level in self-employment, with effort affecting the success probability. For instance, if the worker exerts effort $e\in\{0,1\}$ at a private cost $c>0$, we might have
$\Pr(y_{t}=1\mid\theta,e=1)=\theta\quad\text{and}\quad\Pr(y_{t}=1\mid\theta,e=0)=\phi\,\theta$
for some $\phi\in(0,1)$; exerting effort raises the chance of success, but effort is not observed by the market. This introduces a moral-hazard margin on top of the observability choice.
The qualitative incentives are straightforward. Under $E$, individual output is hidden and wages do not depend on performance in our baseline, so career concerns provide little reason to exert costly effort: a worker who has retreated into hidden employment mainly values the insurance component and, absent additional frictions, will tend to shirk. Under $S$, by contrast, a success is a public signal that improves future wages, while a failure is a public setback. The value of effort in self-employment is therefore tied to how much a marginal success or failure moves the posterior about talent, which in turn depends on the current public state $(\alpha,\beta)$.
This suggests a non-monotone “work–shirk” pattern in reputation space that is closely related to Board and Meyer-ter Vehn, (2013). When the public record is short and beliefs are still diffuse, a single success can substantially raise the posterior mean, so the reputational return to effort in $S$ is high; workers with moderate or promising reputations have strong incentives to work rather than shirk. When the public record is very long and beliefs are precise, each additional success or failure moves $\mu(\alpha,\beta)$ only slightly, so the marginal reputational benefit of effort is smaller; depending on $c$, some high-reputation workers may optimally coast on their existing reputation. At very low reputations, the chance that even a worked-for success would materially improve future wages can be small, so discouraged workers may shirk as well. In short, allowing hidden effort would generate familiar work–shirk dynamics in which effort is most valuable at intermediate reputation levels, echoing the comparative statics of dynamic reputation models under different information structures.
Adding hidden effort would also reinforce the logic of the absorbing region. Once a worker enters the absorbing hidden-employment region, there is little reputational gain from exerting effort in $E$, so both public information production and effort provision are low. By contrast, for types near the cutoff $\hat{\theta}(\alpha,\beta)$, the prospect of escaping absorption by producing a public success can make effort in $S$ particularly attractive. A full solution of the extended model is beyond our scope here, but the discussion above suggests that our main comparative statics—about how risk exposure, transparency, and wage-setting affect the absorbing region and selection into public production—are robust to introducing a hidden effort choice. [^20] We thank an anonymous referee for suggesting that we discuss hidden effort and its connection to work–shirk dynamics in dynamic reputation models, such as Board and Meyer-ter Vehn, (2013).
6.5 Policy implications: taxes and subsidies for public work
The model highlights a simple information externality. When the worker chooses self-employment $S$, the public signal generated in that period not only affects the worker’s own future wages but also improves the allocation of talent in the wider labor market: better information about ability helps firms match workers to jobs more efficiently. By contrast, when the worker chooses hidden employment $E$, individual outcomes remain opaque and the public belief does not move. From a social point of view, workers privately under-produce public signals relative to a planner who internalizes this informational benefit, especially when risk aversion and incomplete insurance make $E$ attractive.
A natural policy lever is to change the relative payoff of $S$ and $E$ through taxes or subsidies on activities that generate public records. Suppose a proportional tax $\tau\in[0,1)$ is levied on self-employment income, so that a success under $S$ yields $(1-\tau)$ instead of $1$. Holding beliefs fixed, this reduces the current-period return to $S$ and raises the talent cutoff $\hat{\theta}(\alpha,\beta)$ in the worker’s indifference condition. In terms of the absorbing hidden-employment region
$\mathcal{A}(\alpha,\beta)=\left\{(\alpha,\beta,\theta):\theta\leq\hat{\theta}(\alpha,\beta)\right\},$
a higher tax expands $\mathcal{A}(\alpha,\beta)$: some types who would have chosen $S$ at a given public state now optimally remain in $E$ and stop generating public information. A symmetric argument applies to a subsidy (a negative $\tau$): by increasing the net payoff from $S$, a subsidy lowers $\hat{\theta}(\alpha,\beta)$, shrinks the absorbing region and induces more experimentation and public signal production. This comparative static is directly analogous to the effects of risk aversion, patience, and wage policies characterized in Proposition 3: policies that effectively make $S$ less attractive (more risky, less remunerative) raise the cutoff and reduce information revelation; policies that make $S$ more attractive do the opposite.
In practice, a planner would trade off the informational benefits of encouraging $S$ against the costs of additional risk and potential misallocation when low-ability workers also experiment. But the qualitative prescription is clear: if the goal is to maximize public information about talent—so that high-$\theta$ workers are identified and deployed efficiently sooner—taxes on activities that generate portable public records (freelancing, open-source contributions, public performance metrics) are counterproductive, while subsidies or insurance for such activities can be welfare-improving. This perspective offers one rationale for policies such as tax breaks for early-career entrepreneurs or support for platforms that document individual performance, and cautions against fiscal measures that inadvertently widen the absorbing region by discouraging public work.
The same logic applies outside labor markets. One can reinterpret the worker as a firm with a privately known product quality and the public signal as a noisy review or rating. Choosing $S$ is then analogous to offering free trials or pay-what-you-want deals that generate public feedback, while choosing $E$ corresponds to selling only through opaque channels where individual customer experiences are not easily observed or shared. High-quality firms are more willing to bear demand risk in order to build a public reputation, whereas low-quality firms prefer the insurance value of opaque pricing or bundled sales.
The absorbing hidden-employment region in our baseline becomes an absorbing hidden-sales region in which firms profitably remain opaque and stop generating public reviews. Comparative statics carry over: environments with better risk sharing, more patient owners, or more transparent technologies induce more firms to experiment with free trials and other reputation-building devices, increasing the public information available about product quality.
@@ 7 Conclusion
Three forces organize the results. First, an insurance–information trade-off: hidden employment offers insurance but suppresses public learning and future outside options. Second, selection: at any public state there is a unique talent cutoff separating workers who continue to self-employ from those who choose hidden employment. Third, pricing sophistication: in the sophisticated regime, wages are set in anticipation of who will apply, so relative to the myopic benchmark they incorporate an adverse-selection discount. Together these forces generate an absorbing hidden-employment region and a pattern in which higher-talent workers select into public production while lower-talent workers select into hidden jobs.
We deliver a simple characterization under both posting regimes, show how the size of the absorbing region responds to risk aversion, patience, and transparency, and show that sophisticated pricing compresses wages, shrinks the absorbing region, and tightens self-selection by shifting the cutoff downward at each public state.
The framework yields transparent empirical implications. Early in a worker’s public track record—when a single public outcome moves beliefs more— workers are more willing to generate public information; once a long public record is in place and additional signals move beliefs less, absorbing hidden employment becomes more prevalent. After public successes (failures), the benchmark predicts wide wage dispersion, whereas sophisticated posting corrects for winner’s-curse effects through application-based inference, compressing the success–failure gap. Across occupations, environments with portable public histories (e.g., open portfolios, platform ratings, contest leaderboards) should feature stronger selection into public production by higher-ability workers and sharper wage responses to public outcomes. [^21] Examples include code and data science (open repositories), academic research (working papers and preprints), design (public portfolios), and online freelancing (platform ratings).
@@
Two design lessons follow. First, organizational transparency is a lever: policies that selectively externalize performance (client-visible tasks, standardized public metrics) can substitute for self-employment as a reputational technology, [^22] For instance, making some client KPIs or review summaries public, adopting external code review, or giving structured, verifiable titles that map to outside‐market signals. reducing the absorbing hidden-employment region and speeding market learning. Second, marketplace curation (platform labels, verified histories) changes the relative returns to public signals and therefore the equilibrium mix of public versus hidden production.
@@
By tying a classic career-concerns logic to environments with coexisting public and hidden performance, the paper provides a compact theory of how selection into self-employment versus hidden jobs and the sophistication of wage setting jointly determine when employment becomes absorbing, how wages reflect reputation, and how much public information about talent different labor markets generate.
Our analysis is also closely related to recent work on dynamic reputation with endogenous records and effort. The stationary, overlapping-generations formulation we adopt follows the spirit of Pei, (2025), who shows how record design and long lifetimes shape incentives in repeated interactions; here, the choice between self-employment and hidden employment plays a similar role in shaping the worker’s public record. The hidden-effort extension discussed above links our selection mechanism to the work–shirk dynamics, highlighting that our main comparative statics about transparency and risk exposure continue to matter once effort provision is added to the observability choice.
Appendix A Proofs
@@
and $U_{E}(\theta;\alpha,\beta)=u(w)+\delta V(\theta;\alpha,\beta)$, the current term implies $\partial U_{S}/\partial\theta>0$; the transition also shifts weight toward $(\alpha{+}1,\beta)$, where $V$ is (weakly) higher by value monotonicity. By contrast, $U_{E}$ depends on $\theta$ only through $V(\theta;\alpha,\beta)$ (same state), hence $dU_{E}/d\theta$ is weakly smaller. Therefore $\Delta(\theta;\cdot,w)$ is strictly increasing and continuous in $\theta$.
$U_{E}$ is strictly increasing and continuous in $w$, while $U_{S}$ is $w$-independent. Hence $\Delta$ is strictly decreasing and continuous in $w$.
@@
Fix a public state $(\alpha,\beta)$ in the absorbing employment region and let $w(\alpha,\beta)$ be the associated wage. For each type $\theta$ let $\Delta(\theta;\alpha,\beta,w(\alpha,\beta))$ denote the expected discounted utility difference between choosing $S$ and choosing $E$ when the current public state is $(\alpha,\beta)$, the wage is $w(\alpha,\beta)$, and future play is given by the equilibrium strategy. By definition of the cutoff $\hat{\theta}(\alpha,\beta)$ we have
$\Delta(\theta;\alpha,\beta,w(\alpha,\beta))\;\leq\;0\quad\text{for all }\theta\leq\hat{\theta}(\alpha,\beta),$
that is, at $(\alpha,\beta)$ these types weakly prefer $E$ to $S$.
Under $E$ the public state is a controlled Markov state with deterministic transition $(\alpha,\beta)\mapsto(\alpha,\beta)$: if the worker chooses $E$ today, the public posterior remains $(\alpha,\beta)$ tomorrow and the firm posts the same wage $w(\alpha,\beta)$ again. Thus, if a type $\theta\leq\hat{\theta}(\alpha,\beta)$ chooses $E$ in the current period, the decision problem he faces next period is identical to the one he faces today: the public state and wage are unchanged and the equilibrium continuation strategy profile is the same.
Let $V_{t}(\alpha,\beta,\theta)$ denote the worker’s continuation value at the start of period $t$, and suppose that at some date $t_{0}$ the public state is $(\alpha,\beta)$ and $\theta\leq\hat{\theta}(\alpha,\beta)$. Consider the stationary strategy that prescribes $E$ in every period $t\geq t_{0}$. Under this strategy the public state remains $(\alpha,\beta)$ forever, so the Bellman equation at $t_{0}$ can be written as
$V_{t_{0}}(\alpha,\beta,\theta)=\max\{U_{S}(\alpha,\beta,\theta),\,U_{E}(\alpha,\beta,\theta)\},$
where $U_{S}$ and $U_{E}$ denote the current-period payoff plus discounted continuation value from playing $S$ and $E$, respectively, given the equilibrium continuation strategy. By construction, $U_{S}(\alpha,\beta,\theta)-U_{E}(\alpha,\beta,\theta)=\Delta(\theta;\alpha,\beta,w(\alpha,\beta))\leq 0$, so the maximum is attained at $E$:
$V_{t_{0}}(\alpha,\beta,\theta)=U_{E}(\alpha,\beta,\theta).$
Exactly the same argument applies at $t_{0}+1,t_{0}+2,\ldots$ because the public state and wage remain $(\alpha,\beta)$ and the continuation strategy is stationary. Hence at every date $t\geq t_{0}$ there is no profitable one-period deviation from the strategy that prescribes $E$.
By the one-step deviation principle for stationary dynamic programming problems, the strategy that chooses $E$ in every period $t\geq t_{0}$ is sequentially optimal for types $\theta\leq\hat{\theta}(\alpha,\beta)$ starting from $(\alpha,\beta)$. In particular, once such a type enters the absorbing employment region and chooses $E$, he will optimally continue to choose $E$ forever and never produce public signals again, so the region is absorbing. ∎
@@
For any $w\in(0,1)$, $\Delta(0;\cdot,w)<0$ (since $U_{E}(0)=u(w)>0=U_{S}(0)$) and $\Delta(1;\cdot,w)>0$ (since $U_{S}(1)=1+\delta V(\cdot)\!>\!u(w)+\delta V(\cdot)=U_{E}(1)$).
By continuity and strict single crossing in $\theta$ there exists a unique $\phi(w)\in(0,1)$ with $\Delta(\phi(w);\cdot,w)=0$, and $\phi(w)$ is strictly decreasing in $w$ by the implicit function theorem and Lemma 1(2).
[…]